OpenStudy (smartpanda1):

How do you do 19-6(-k-2) and why do you do it?

OpenStudy (3mar):

This is how: \[19-6(-k-2)=19-6*-k-6*-2=19+6k+12=(19+12)+6k=31+6k\] but for why: because you need it

OpenStudy (hoot_hoot):

(19 + 12) + 6k = 31 + 6k

OpenStudy (hoot_hoot):

19 - 6(-k -2) = 19 +6k + 12 (DPMA) = 31 + 6k (Adding Like Terms)